A stock with a current price of $300, can go either up by 20%, or down by 15%. You have the opportunity to buy an American call option on this stock for $65, which matures at the end of one period, with an exercise price of 275.

a-      Please draw the trees of the stock and option values

b-      What is the price of the call option today if the risk free rate is 8%.

c-      What is your decision for this investment opportunity? Explain your answer


1
Expert's answer
2022-01-10T05:12:29-0500

Solution:

Given:

Spot price, "S =300"

"\begin{aligned} u &=1.20 \\ d &=0.85 \end{aligned}"

American call option

Two period.

Exercise, x=275

Risk free rate, r=8%

Solution:

Risk neutral probability "=\frac{(1+r)-d}{u-d}"

"\begin{aligned}

&=\frac{(1+8 \%)-0.85}{1.20-0.85}=0.6571 \\

&up\ move=65.71\% \\

&down\ move=34.29 \%

\end{aligned}"

"\begin{aligned}

\text { up move } &=62.22 \% \\

\text { down move } &=37.78 \%

\end{aligned}"

(a) Binomial tree for american call:


"\begin{gathered}
120.36=\frac{(193.75 \times 62.22 \%)+(25 \times 37.78 \%)}{(1+8 \%)} \\
v_{S} \\
375-275=100
\end{gathered}"

So, we select 120.36

(2)

"\begin{aligned}

14 . 40 &=\frac{(25 \times 62.22 \%)}{(1+8 \%)} \\

& \text { vs } \\

240-275 &=\text { out of the money }

\end{aligned}"

Se, we select 14.40.

(3)

"\begin{gathered}

74.38=\frac{(120.36 \times 62.22 \%)+(14.40 \times 37.78 \%)}{(1+8 \%)} \\

\text { VS } \\

300-275=25

\end{gathered}"

So we select 74.38.

(b) Price of the call option =74.38


(c) Given American coll option price =$ 65

Calculated American calloption price = $74.38

As given American call option price and calculated American call option price are diffecent, there is a scope of profit from this investment.


Buy the American call option al lower of the above prices.